Formaldehyde is manufactured by the catalytic oxid
Answer #1

First consider the reaction (ii)

HCHO + 1/2 O2 = HCOOH ...... (ii)

from stiochiomentry 1 mol of HCHO + 1/2 mol of O2 = 1 mol of HCOOH

Product composition contains 0.6 mol of HCOOH,

hence from stoichiometry, Total no. moles of HCHO reacted (in reaction ii) to form HCOOH are: 0.6 mol ....... (a)

& Total no. of moles of O2 reacted (in reaction ii) to form HCOOH are: 0.3 mol ................................................(b)

Total no. moles of HCHO formed before the second reaction are:moles of HCHO in product + moles of HCHO consumed to form HCOOH = 3.1 +0.6 = 3.7 moles.

From reaction (i)

CH3OH + 1/2 O2 = HCHO + H2O

From stoichimetry: 1 mol CH3OH + 1/2 mol O2 = 1 mol HCHO + 1 mol H2O

Total no. of moles of of HCHO formed are : 3.7 mol

Hence from stoichiometry, total no. moles of CH3OH reacted to form HCHO are: 3.7 mol .......................................(c)

& total no. moles of O2 reacted to form HCHO are: 1.85 mol ...................................................................................(d)

Now, total no. O2 consumed in reaction (i) & (ii) are: 1.85 + 0.3 = 2.15 ...................................................................(e)

a) % conversion of methanol to formaldehyde is given by:

% Conversion = (moles of methanol reacted to form HCHO / Total no. moles of methanol in the feed)

=(3.7 / (3.7+8.6)) * 100 = (3.7/12.3)*100 = 30.08

% Conversion of methanol to HCHO is 30.08

b) % Conversion of methanol to HCOOH:

Two reaction:   CH3OH + 1/2 O2 = HCHO + H2O

  HCHO + 1/2 O2 = HCOOH

Overall reaction CH3OH + O2 = HCOOH + H2O

Total no. moles of HCOOH formed: 0.6 mol

Hence Total no, moles of methanol reacted to form HCOOH is: 0.6 mol (From stoichiometry)

% Conversion of methanol to HCOOH= (moles of CH3OH reacted to form HCOOH/ total no. moles of CH3OH in feed)

= (0.6/12.3) * 100 = 4.88

% Conversion of methanol to HCOOH is 4.88 %

C) Molar ratio of Air to Methanol

Total no. of moles of methanol in feed= 12.3 mol

Total no. of mol of N2 in product = 68 mol

So, Total no. of moles of air in feed = (68/.79) = 86.08 mol (Assume: Air compostion as 21% O2 & 79 % N2)

So, Air to methanol ratio= 86.08 / 12.3 = 6.998 = 7

d) Yeild of Formeldehyde

Yeild = (Moles of HCHO Formed) /(Total moles of Methanol in feed - Total moles of Methanol in product)

= 3.1 / (12.3-8.6) = 0.8378

Selectivity of Formaldehyde

Selectivity = moles of desired product formed (HCHO) / moles of undesired product formed (HCOOH)

= 3.1 / 0.6 = 5.17